Friday, 21 August

Special Activities

\underline{14:00 – 18:00 H 104}

Mathematical Software Chair: Winfried Neun, ZIB, Berlin, Germany In this session, a broad spectrum of mathematical software including symbolic and numerical systems as well as interfaces and problem solutions will be presented. Details of the presentations can be found on the ICM’98 web server at URL http://elib.zib.de/ICM98/MathSoftware .

14:00 – 14:30 John H. Maddocks and Randy C. Paffenroth, EPF Lausanne, Switzerland VBM

14:40 – 15:10 Horst G. Zimmer et. al., Universität Saarbrücken, Germany SIMATH

15:15 – 15:45 Karim Belabas, Université Paris–Sud, Orsay, France PARI / GP

15:45 – 16:15 Break

16:15 – 16:45 Christian Wieners et. al., Universität Stuttgart, Germany UG

16:50 – 17:20 Elias Deeba, University of Houston–Downtown, Texas, USA Learning Linear Algebra Using an Interactive Text

17:30 – 18:00 David M. Gay and Robert Fourer, Murray Hill, N.J., USA AMPL \underline{19:30 – 21:00 H 104}

Events and policies: Effects on women in mathematics Panel discussion organized by women from the Association for Women in Mathematics (AWM), the European Women in Mathematics (EWM), and the Committee on Women and Mathematics of the European Mathematical Society

\underline{21:15 – 22:00 H 104}

Women and mathematics across cultures Film, directed by Marjatta Näätänen (Helsinki, Finland) in collaboration with Bodil Branner (Lyngby, Denmark), Kari Hag (Trondheim, Norway), and Caroline Series (Warwick, UK)